What is compressibility factor (z) and it`s formula

Grade:11

1 Answers

Arun
6 years ago

The compressibility factor (Z), also known as the compression factor or the gas deviation factor, is a correction factor which describes the deviation of a real gas from ideal gas behavior. It is simply defined as the ratio of the molar volume of a gas to the molar volume of an ideal gas at the same temperature and pressure

In statistical mechanics the description is:

{\displaystyle Z={\frac {pV}{nRT}}}{\displaystyle Z={\frac {pV}{nRT}}}

where p is the pressure, n is the number of molea of gas, {\displaystyle T}T is the absolute temperature, and {\displaystyle R}R is the gas constant